A beam of parallel light is incident normally on a diffraction grating having 550 lines per mm. A

 telescope is used to observe the second order in the spectrum. Calculate the angular separation in radians of two spectra lines of wavelengths 559nm and 563nm.


1
Expert's answer
2021-03-14T19:47:30-0400

Here,

"\lambda=559\ nm\\\lambda+\Delta \lambda=563\ nm"

"\Delta \lambda=(563-559)nm=4\ nm"

"m=1\\d=1/550\ =2\times 10^{-6}m"

 Now,

sinα1= mλ1/d

sinα2= mλ2/d

Substracting the two gives:

sinα1- sinα2= m(λ1-λ2)/d

Since the angles are small, we approximate sinα≈α to get the separation:

Angular separation "\alpha_1- \alpha_2=\dfrac{m(\lambda_1-\lambda_2)}{d}"

"= \dfrac{1(563-559)\times 10^{-9}} {2\times 10^{-6}}=2\times 10^{-3}"